OpenStudy (anonymous):

i dont know this one... integrate arctan (x/3) dx

OpenStudy (anonymous):

Think you can say x = Tan y = sin y/cosy and use the f'/f integration.

OpenStudy (anonymous):

Or x/3 = Tan y maybe.

OpenStudy (anonymous):

Or maybe the trick with multiplying by 1.

OpenStudy (anonymous):

i am working with arc tan

OpenStudy (anonymous):

a little advanced...are you at integration by parts yet?

OpenStudy (anonymous):

Firstly, I would let p=x/3 dp=(1/3)dx 3dp=dx This makes your integral: \[3 \int\limits \arctan(p)dp\] Now, you must use integration by parts. You let u=arctan(p) (something easy to differentiate) and dp=dv (something easy to integrate) So: \[du=\frac{dp}{1+p^2}\] And v=p Now use the formula: \[uv-\int\limits v du \rightarrow 3p \arctan(p)-3\int\limits \frac{p}{p^2+1}dp\] For the second integral use the substitution s=p^2+1 => ds=2p dp => (1/2)ds=dp Your integral now becomes: \[3p \arctan(p)-\frac{3}{2} \int\limits \frac{ds}{s}=3p \arctan(p)-\frac{3}{2}\ln|s|+C\] Plugging back in: \[3p \arctan(p)-\frac{3}{2}\ln|p^2+1|+C\] Then p. \[x\arctan(\frac{x}{3})-\frac{3}{2}\ln|(\frac{x}{3})^2+1|+C\] If you look it up on wolfram they factor out a stuff inside the ln and let it be absorbed by C, that is not really necessary imo.

OpenStudy (anonymous):

ok,, perfect.. let me try it...